# CurbNet: Curb Detection Framework Based on LiDAR Point Cloud Segmentation

### v1

@@ IV-F Real Scene Experiment
The curb detection results in the 5 scenes are illustrated in Fig. 10, where we selected 8 representative images from each scene for visual analysis. Overall, our method achieved commendable curb detection results in each scene, further demonstrating the robust curb feature extraction capability of the CurbNet model. In individual cases, our proposed method accurately detected not only the evident, extensive curbs, but also achieved remarkable results on small-scale curbs, which are typically less conspicuous and easily overlooked, as seen in Scene B (image d) and Scene D (images b and c). This success can be attributed to the model’s design focusing on multi-scale feature fusion and multi-channel feature extraction. Similarly, our method also exhibited superior detection performance in complex intersection scenarios with irregular curb distributions (Scene B and Scene C). Particularly in Scene B, the method precisely detected all curbs present in the LiDAR point cloud, a feat attributable to the powerful curb feature extraction ability of the MSCA module.
Finally, this paper also quantitatively evaluates the real scene experiments by testing key metrics, the results as shown in Table. VI. By comparing with manually annotated ground truth data, CurbNet achieved an average Precision, Recall, and F1 score of 0.8462, 0.8443, and 0.8453, respectively, across the five scenes. Among them, the Recall and F-1 score indicators obtained in Scene A are the highest, which are 0.8821 and 0.8569 respectively, and the Precision indicators obtained in Scene B are the highest 0.8950. These results corroborate with those obtained from dataset testing, further substantiating the excellent detection performance and generalizability of the CurbNet model.
@@ VI Conclusion
In this paper, we established the 3D-Curb dataset, comprising 7,100 frames. To our knowledge, this is currently the largest and most diverse curb point cloud dataset with the most extensive range of annotated categories. Notably, this is also the first dataset to feature 3D point cloud annotations for curbs, which will significantly aid future related research. Within the CurbNet framework, we introduced the Multi-Scale and Channel Attention (MSCA) module, addressing the challenges of uneven distribution of curb features on the xy-plane and the reliance on high-frequency z-axis features. Additionally, we developed an adaptive weighted loss function group to resolve the imbalance in the number of curb point clouds relative to other categories. Extensive experiments on both the NRS and 3D-Curb datasets demonstrated that our approach outperforms the current leading curb detection and point cloud segmentation models. In tolerance experiments, CurbNet achieved over 0.95 average performance in Precision, Recall, and F-1 score metrics at just 0.15m tolerance, setting a new standard. Furthermore, our post-processing approach of multi-clustering and curve fitting effectively eliminated noise in the curb results, enhancing the Precision, Recall, and F-1 score metrics to 0.8744, 0.8648, and 0.8696, respectively. Finally, the excellent detection performance and generalization of our proposed method were further verified in real scene experiments.

### v2

@@ IV-F Real Scene Experiment
The curb detection results in the 5 scenes are illustrated in Fig. 11, where we selected 8 representative images from each scene for visual analysis. Overall, our method achieved commendable curb detection results in each scene, further demonstrating the robust curb feature extraction capability of the CurbNet model. In individual cases, our proposed method accurately detected not only the evident, extensive curbs, but also achieved remarkable results on small-scale curbs, which are typically less conspicuous and easily overlooked, as seen in Scene B (image d) and Scene D (images b and c). This success can be attributed to the model’s design focusing on multi-scale feature fusion and height channel feature extraction. Similarly, our method also exhibited superior detection performance in complex intersection scenarios with irregular curb distributions (Scene B and Scene C). Particularly in Scene B, the method precisely detected all curbs present in the LiDAR point cloud.
Finally, we also quantitatively evaluate the real scene experiments by testing key metrics, as shown in Table. VI. By comparing with manually annotated ground truth, CurbNet achieved an average Precision, Recall, and F1 score of 0.8462, 0.8443, and 0.8453, respectively, across the five scenes. Among them, the Recall and F-1 score indicators obtained in Scene A are the highest, which are 0.8821 and 0.8569 respectively, and the Precision indicator obtained in Scene B is the highest 0.8950. These results corroborate with those obtained from dataset testing, further substantiating the excellent performance and generalizability of the CurbNet.
@@
IV-G Time Consumption
In practical applications for intelligent vehicles, real-time curb detection is critical. As shown in Table. VII, we have evaluated the time consumption for both the model inference and post-processing stages of our curb detection framework. To ensure the efficiency of the entire processing framework, the model inference is executed on the GPU, while the post-processing runs on the CPU.
In various scenarios, since the LiDAR used and the number of input point clouds are different, the time consumption of computational processing is also different. Nevertheless, the CurbNet framework consistently achieves an overall real-time performance exceeding 15 Hz during both model inference and post-processing stages. Notably, the post-processing stage operates faster than the model inference stage, ensuring the smooth and efficient operation of the entire framework.
TABLE VII: Comparison of time consumption in model inference and post-processing, the unit is (FPS/ms).
Stage 3D-Curb NRS Real Scene
Model Inference 20.69 / 48.33 18.18 / 55.01 16.22 / 61.65
Post-Processing 26.96 / 37.09 26.08 / 38.34 19.02 / 52.58
@@ VI Conclusion
In this paper, we established the 3D-Curb dataset, comprising 7,100 frames. To our knowledge, this is currently the largest and most diverse curb point cloud dataset with the most extensive range of annotated categories. Notably, this is also the first dataset to feature 3D point cloud annotations for curbs, which will significantly aid future related research. Within the CurbNet framework, we introduced the Multi-Scale and Channel Attention (MSCA) module, addressing the challenges of uneven distribution of curb features and the reliance on high-frequency z-axis features. Additionally, we introduce a novel adaptive loss function group to resolve the imbalance in the number of curb point clouds relative to other categories. Extensive experiments on both the NRS and 3D-Curb datasets demonstrated that our approach outperforms the current leading curb detection and point cloud segmentation models. In the tolerance experiments, CurbNet achieved over 0.95 average performance in Precision, Recall, and F-1 score metrics at just 0.15m tolerance, setting a new standard. Furthermore, our post-processing approach of multi-clustering and curve fitting effectively eliminated noise in the curb results, enhancing the Precision, Recall, and F-1 score metrics to 0.8744, 0.8648, and 0.8696, respectively. Finally, the excellent detection performance and generalization of our proposed method were further verified in real scene experiments.
@@
Acknowledgments
This work was supported by Guangdong Basic and Applied Basic Research Foundation (No. 2021B1515120032), Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0618).